OpenStudy (anonymous):

The tip of a tuning fork goes through 440 complete vibrations in a time of 0.510s. Find the angular frequency and the period of the motion.

OpenStudy (schrodingers_cat):

f = 440/.510 = 862.7 hz Angular Frequency, w = 2(pi)862.7 = 5420 T = 1/(862.7) = .00116 s I think this what your were looking for hope this helps :)

OpenStudy (anonymous):

you're awesome! Thanks so much!

OpenStudy (masumanwar):

frequency and then angular frequency 2